# Machine learning-driven conservative-to-primitive conversion in hybrid piecewise polytropic and tabulated equations of state

### v2

@@ 4 Conclusions
This work introduces a novel ML-driven method for accelerating C2P inversions in relativistic hydrodynamics simulations, with a focus on hybrid piecewise polytropic and tabulated equations of state. By employing feedforward neural networks optimized with TensorRT, we achieve substantial performance improvements over traditional CPU solvers, offering a compelling alternative to computationally expensive iterative methods, while maintaining high accuracy. Our results demonstrate that the TensorRT-optimized neural networks can process large datasets significantly faster, achieving up to 25 times the inference speed of traditional methods.
@@
Future work will explore several key directions to refine and expand this approach. First, adapting the models to handle a broader range of equations of state will improve the versatility of this method across different simulation contexts. Second, exploring alternative network architectures could further enhance both accuracy and inference speed. Additionally, continued optimization of TensorRT, including advanced parallelization strategies and scaling across multiple GPUs, promises even greater reductions in computational time, enabling simulations of larger and more complex astrophysical systems. These improvements will be critical for advancing high-resolution simulations in numerical relativistic hydrodynamics.
@@
We believe that ML-driven methods, particularly those incorporating TensorRT optimization, will play an essential role in advancing the field of general relativistic hydrodynamics and numerical relativity more broadly. To facilitate further validation and extension of these findings, we have made the software developed for this study publicly available at the following GitHub repository [35].
5 Acknowledgements
E.A.H. gratefully acknowledges support from National Science Foundation award OAC-1931561 and OAC-2209892. R.H. gratefully acknowledges support from Natural Science Foundation award OAC-2103680, OAC-2004879, OAC-2310548, OAC-2005572, and OAC-2411068. This research used the Delta advanced computing and data resource which is supported by the National Science Foundation (award OAC 2005572) through ACCESS-CI [36] allocation PHY160053. Delta is a joint effort of the University of Illinois Urbana-Champaign and its National Center for Supercomputing Applications. This research used the DeltaAI advanced computing and data resource, which is supported by the National Science Foundation (award OAC 2320345) and the State of Illinois. DeltaAI is a joint effort of the University of Illinois Urbana-Champaign and its National Center for Supercomputing Applications. We further acknowledge the use of Matplotlib [37] and Seaborn [38] for the generation of figures in this publication.

### v3

@@ IV Conclusions
This work introduces a novel ML-driven method for accelerating C2P inversions in relativistic hydrodynamics simulations, with a focus on hybrid piecewise polytropic and tabulated equations of state. By employing feedforward neural networks optimized with TensorRT, we achieve substantial performance improvements over traditional CPU solvers, offering a compelling alternative to computationally expensive iterative methods while maintaining high accuracy. Our results demonstrate that the TensorRT-optimized neural networks can process large datasets significantly faster, achieving up to 25 times the inference speed of traditional methods. The success of this approach is rooted in the neural network’s ability to efficiently learn and represent the inherent symmetries and complex functional relationships within the EOS, effectively creating a direct mapping that bypasses iterative numerical solvers.
@@
Future work will explore several key directions to refine and expand this approach. First, adapting the models to handle a broader range of equations of state will improve the versatility of this method across different simulation contexts. Second, exploring alternative network architectures, such as those incorporating physics-informed layers or adaptive activation functions to better handle physical discontinuities like phase transitions, could further enhance both accuracy and inference speed. Third, the models must be extended to handle full three-dimensional velocities to be fully integrated into production-level GRMHD codes. Additionally, continued optimization of TensorRT, including advanced parallelization strategies and scaling across multiple GPUs, and careful exploration of lower-precision formats like INT8, potentially with quantization-aware training, promises even greater reductions in computational time, enabling simulations of larger and more complex astrophysical systems. These improvements will be critical for advancing high-resolution simulations in numerical relativistic hydrodynamics.
@@
We believe that ML-driven methods, particularly those incorporating TensorRT optimization, will play an essential role in advancing the field of general relativistic hydrodynamics and numerical relativity more broadly. To facilitate further validation and extension of these findings, we have made the software developed for this study publicly available at: https://github.com/semihkacmaz/C2PNets (accessed on 27 August 2025).
Appendix A Model Architecture Exploration and Training History
In this study, we explored a wide range of multi-layer perceptron (MLP) architectures to identify models that offer an optimal balance between predictive accuracy and inference speed. The models presented in the main text—NNC2PS, NNC2PL, and NNC2P_Tabulated—were the result of this systematic exploration.
Our findings, summarized in Table 3, demonstrate a clear “sweet spot” for model complexity. Architectures smaller than our chosen models (e.g., NNC2PS-Tiny) offered lower parameter counts but with a notable drop in accuracy. Conversely, models that were significantly wider or deeper than our selections provided only marginal accuracy gains for a substantial increase in parameter count and computational cost (e.g., NNC2PL-Wide, NNC2PL-Deep-7). This trend of diminishing returns is evident across both EOS types.
Notably, excessively deep architectures (e.g., the 10- and 13-layer models) consistently exhibited training instability or yielded worse performance, reinforcing our choice of moderately sized networks as the most effective and efficient solution for this regression task.
To demonstrate the stability of our training procedure, Figures 5 and 6 show the training and validation loss curves for the final three models used in this work. The curves illustrate smooth convergence to a low loss value with no signs of significant overfitting, validating our training methodology.
Table 3: Explored architectures and validation accuracy ($L_{1}$ error) for both EOS models. The selected models are shown in bold. The validation error measures the model’s performance on unseen data.
Model Name Hidden Layers (Neurons per Layer) Total Parameters Validation $L_{1}$ Error
Piecewise Polytropic EOS
NNC2PS-Tiny [300, 100] ~31 k $5.8\times 10^{-7}$
NNC2PS-Shallow [800] ~3 k $6.5\times 10^{-7}$
NNC2PS [600, 200] ~123 k $\mathbf{4.5\times 10^{-7}}$
NNC2PS-Wide [800, 400] ~324 k $4.1\times 10^{-7}$
NNC2PL-Small [512, 256, 128, 64] ~180 k $3.9\times 10^{-7}$
NNC2PL-Medium [1024, 512, 256, 128] ~690 k $3.2\times 10^{-7}$
NNC2PL [1024, 512, 256, 128, 64] ~707 k $\mathbf{2.8\times 10^{-7}}$
NNC2PL-Wide [2048, 1024, 512, 256, 128] ~2.8 M $2.5\times 10^{-7}$
NNC2PL-Deep-7 [1024, 1024, 512, 512, 256, 128, 64] ~2.4 M $2.9\times 10^{-7}$
NNC2PL-Deep-10 [1024, 1024, 512, 512, 256, 256, 128, 128, 64, 64] ~3.5 M $3.1\times 10^{-7}$
NNC2PL-SuperDeep 13 Layers ~5 M Failed to Converge
Tabulated EOS (LS220)
NNC2P_Tab-Tiny [512, 256, 128] ~165 k $9.5\times 10^{-3}$
NNC2P_Tab-Small [1024, 512, 256, 128] ~690 k $8.8\times 10^{-3}$
NNC2P_Tabulated [1024, 512, 256, 128, 64] ~707 k $\mathbf{8.0\times 10^{-3}}$
NNC2P_Tab-Wide [2048, 1024, 512, 256, 128] ~2.8 M $7.7\times 10^{-3}$
NNC2P_Tab-Deep-7 [1024, 1024, 512, 512, 256, 128, 64] ~2.4 M $8.2\times 10^{-3}$
NNC2P_Tab-Deep-10 [1024, 1024, 512, 512, 256, 256, 128, 128, 64, 64] ~3.5 M $8.5\times 10^{-3}$
NNC2P_Tab-SuperDeep 13 Layers ~5 M Failed to Converge
(a) NNC2PS
(b) NNC2PL
Figure 5: Training and validation loss curves for the piecewise polytropic EOS models. The smooth convergence demonstrates a stable training process for (a) NNC2PS and (b) NNC2PL.
Figure 6: Training and validation loss curves for the tabulated EOS model, NNC2P_Tabulated. The smooth convergence demonstrates a stable training process without significant overfitting.